john and tom are brothers. Tom is seven years older than john. In four years tom will be twice as old as john. how old is tom? how old is john

Answer by mananth(16946)

john is x years
Tom is x+7 years

4 years hence
John -------- x+4
Tom ---------x+7+4 =x+11
x+11=2(x+4)
x+11=2x+8
2x-x=11-8
x=3 that's john's age
Tom'a age = 3+7 =10 years